### v3

# Oganesson versus Uranium Hydrogen-like Ions from the Viewpoint of Old Quantum Mechanics

Abstract.
We compare, within the framework of the old Bohr-Sommerfeld atomic model, Uranium (Z=92)(Z=92)\/ versus hypothetical Oganesson (Z=118)(Z=118)\/ relativistic hydrogen-like ions and beyond. Existence of a self-intersecting orbit in the super strong static Coulomb field is demonstrated with the aid of Mathematica computer algebra system. A possibility of a similar ‘Oganesson-type’ effect in a strong gravitational field is also mentioned. Appearance of a multi self-intersecting trajectory is demonstrated in a hypothetical Unbibium, an element with Z = 122.
Key words and phrases:
Old quantum mechanics, Sommerfeld fine structure formula, transuranium elements, Uranium, Oganesson, Unbibium, Mathematica computer algebra system.
1991 Mathematics Subject Classification:
Primary 81-01, 81-03. Secondary 81C
Hence it was necessary to carry out the calculation as far as possible according to the method of the older quantum theory; the inferences drawn can then later be taken over directly into wave mechanics.
Arnold Sommerfeld [30, p. 258]
1. Introduction: Fine Structure Formula and Orbits
The original Sommerfeld fine structure formula is given by [29, 30]:
$\frac{E_{n_{r},n_{\theta}}}{mc^{2}}=\left[1+\frac{\alpha^{2}Z^{2}}{\left(n_{r}+\left(n_{\theta}^{2}-\alpha^{2}Z^{2}\right)^{1/2}\right)^{2}}\right]^{-1/2}\qquad(Z<137),$ (1.1)
where $n_{r}$ (the radial quantum number) and $n_{\theta}$ (the azimuthal quantum number) are positive integers. This result made it possible to explain, for the first time, the fine structure of spectral lines. (For further details, see [2, 3, 8, 17, 25] and the references therein.)
In standard geometrical terms, the orbit equation is
$\frac{1}{r}=\frac{1+\epsilon\cos\left(\omega\theta\right)}{a\left(1-\epsilon^{2}\right)}.$ (1.2)
With the eccentricity $\epsilon$, we have, for $\omega\theta=\phi=0$, the perihelion distance $r_{\text{min}}=a(1-\epsilon)$, and for $\omega\theta=\phi=\pi$, the aphelion distance $r_{\text{max}}=a(1+\epsilon)$.
As stated in original works [5, 18, 30], classical relativistic Kepler orbits have the form of conic sections as in the nonrelativistic case, but with a new angular variable $\phi=\omega\theta.$ Thus, for elliptical orbits (bound states), the motion from one perihelion ($\phi=0$) to the next ($\phi=2\pi$) requires $\theta=2\pi/\omega\,$, with a per-revolution shift of $\Delta\theta=2\pi/\omega-2\pi.$ [^1] We follow the terminology of classical work [30] where sufficiently small $\Delta\theta$ are discussed.
Quantized values of parameters of the electron’s elliptical orbits (1.2) are evaluated in [3] as follows
$\omega_{n_{\theta}}n_{\theta}$ $=$ $\left(n_{\theta}^{2}-\alpha^{2}Z^{2}\right)^{1/2},$ (1.3)
$\epsilon_{n_{r},n_{\theta}}$ $=$ $\sqrt{n_{r}}\cdot\frac{\left(n_{r}+2\sqrt{n_{\theta}^{2}-\alpha^{2}Z^{2}}\right)^{1/2}}{n_{r}+\sqrt{n_{\theta}^{2}-\alpha^{2}Z^{2}}},$ (1.4)
$a_{n_{r},n_{\theta}}$ $=$ $\frac{a_{0}}{Z}\left(n_{r}+\sqrt{n_{\theta}^{2}-\alpha^{2}Z^{2}}\right)$
$\times\sqrt{\alpha^{2}Z^{2}+\left(n_{r}+\sqrt{n_{\theta}^{2}-\alpha^{2}Z^{2}}\right)^{2}},$
where $a_{0}=\hbar^{2}/(me^{2})$ is the familiar Bohr radius and $\alpha=e^{2}/(\hbar c)$ is the fine structure formula. These formulas generalize the circular orbits. [^2] Classical solutions of the relativistic Kepler problem are also discussed in [11, pp. 481–482] and [18, pp. 100–102].
2. Some Trajectories of Finite Motion: Uranium vs Oganesson and Beyong
We believe that utilizing the Bohr–Sommerfeld model to contrast $U^{91+}$, $Og^{117+}$, and even hypothetical $Ubb^{121+}$ hydrogen-like ions gives a fresh pedagogical and historical perspective – it links early quantum theory [29, 30] to modern superheavy transuranium relativistic systems [22] in the quasi-classical approximation (using the Dirac equation as a next step [3]).
2.1. Uranium
Uranium (U) is a naturally occurring, radioactive, heavy metal with atomic number Z=92Z=92\/, known for its ability to release concentrated energy through nuclear fission, which powers nuclear reactors and weapons. Quantum electrodynamics of the hydrogen-like Uranium ion had been studied in [12, 13] (see also [10, 19, 27, 28, 31, 22] and references therein). We use Mathematica for the following animations [4].
In the Bohr-Sommerfeld atomic model under consideration, even for the strong electric field of the Uranium nucleus, a single electron stays on classical relativistic elliptical orbits with advancing perihelion, in a form of open “rosettes” [5, 18, 30] (see Figure 1 with the standard orbit ‘winding number’ one).
Figure 1. Kepler’s Elliptical Counterclockwise Motion in Uranium Hydrogen-like Ion $U^{91+}\/.$
In the above animation of the relativistic Kepler motion of an electron in hydrogen-like Uranium ion, $U^{91+}$, when $Z=92$, the nucleus is situated in the fixed focus at the origin. For this quantum state, we choose $n_{r}=n_{\theta}=1$ in the fine structure formula (1.1). By (1.3)–(1), one gets: $\omega=0.741\ 135$, $\epsilon=0.904\ 882$, and a=0.035 3163a=0.035\ 3163, in Bohr’s atomic units. The perihelion and aphelion move along two concentric circles around the nucleus with radii: $r_{\text{min}}=a(1-\epsilon)=0.003\ 3592\ 09$ and $r_{\text{max}}=a(1+\epsilon)=0.067\ 2734\ 72$, respectively. In the animation, the geometrical loci of the successive perihelia and aphelia, the outer and inner envelopes of the orbit, are not shown for simplicity. Counterclockwise rotation of the ellipse over one period is given by $\Delta\theta=2\pi\left(1/\omega-1\right)=$ 2.194 61.2.194\ 61.
2.2. Oganesson
Oganesson (Og) is the heaviest synthetic element on the periodic table, with atomic number Z=118Z=118\/. Discovered by scientists at the Joint Institute for Nuclear Research in Russia and Lawrence Livermore National Laboratory in 20062006\/ [22], it is named after physicist Yuri Oganessian (see [7, 9, 14, 16, 20, 21] and references therein).
Our Mathematica animations show that in a super strong electric field of the Oganesson nucleus, the classical relativistic finite Kepler orbits of a single electron may change its topology (see Figure 2 with the orbit ‘winding number’ two [^3] The so-called “chemical orbit” or “double necklace” in terminology of the American cosmologist John Archibald Wheeler [6, 24, 32]. ).
Figure 2. Kepler’s ‘Elliptical’ Clockwise Motion with Self-intersection in Hypothetical Oganesson Hydrogen-like Ion $Og^{117+}\/.$
In the above animation of the relativistic Kepler motion of an electron in hydrogen-like Oganesson ion, $Og^{117+}$, when $Z=118$, the nucleus is situated in the fixed focus at the origin. For this quantum state, we choose $n_{r}=n_{\theta}=1$ in the fine structure formula (1.1). By (1.3)–(1), one gets: $\omega=0.508\ 457$, $\epsilon=0.941\ 479$, and a=0.022 2041a=0.022\ 2041, in Bohr’s atomic units. The perihelion and aphelion move along two concentric circles around the nucleus with radii: $r_{\text{min}}=a(1-\epsilon)=0.001\ 2994$ and $r_{\text{max}}=a(1+\epsilon)=0.043\ 1087$, respectively. In the animation, the geometrical loci of the successive perihelia and aphelia, the outer and inner envelopes of the orbit are not shown for simplicity. Total rotation of the self-intersecting ‘ellipse’ over one period is given by $\Delta\theta=2\pi\left(1/\omega-1\right)=6.074\ 18.$
[…]
Figure 3. Kepler’s ‘Elliptical’ Counterclockwise Motion with Four Self-intersections in Hypothetical Unbibium Hydrogen-like Ion $Ubb^{121+}\/.$
2.3. Unbibium
As another example, let us discuss a hypothetical element with $Z=122$, which has a temporary name, Unbibium (Ubb), one of a yet-to-be-discovered superheavy element at the end of periodic table.
In the above animation of the relativistic Kepler motion of an electron in a hypothetical hydrogen-like Unbibium ion, $Ubb^{121+}$, when $Z=122$, the nucleus is situated in the fixed focus at the origin. For this quantum state, we choose $n_{r}=n_{\theta}=1$ in the fine structure formula (1.1). By (1.3)–(1), one gets: $\omega=0.455\ 419$, $\epsilon=0.949\ 782$, and a=0.020 3534a=0.020\ 3534, in Bohr’s atomic units.
The perihelion and aphelion move along two concentric circles around the nucleus with radii: $r_{\text{min}}=a(1-\epsilon)=0.001\ 02211$ and $r_{\text{max}}=a(1+\epsilon)=0.039\ 6847$, respectively. In the animation, the geometrical loci of the successive perihelia and aphelia, the outer and inner envelopes of the orbit, see (see Figure 3, are not shown for simplicity). Rotation of the ellipse over one period is given by $\Delta\theta=2\pi(1/\omega-1)=7.513\ 3.\/$ The ‘winding’ number is three.
At the point of the first self-intersection one gets, approximately, r(10.002448573515926)=r(3.7940322175405243)=0.00234067r(10.002448573515926)=r(3.7940322175405243)=0.00234067 with r(10.002448573515926)r(3.7940322175405243)=0.r(10.002448573515926)-r(3.7940322175405243)=0.\/ Here, $\omega(10.002448573515926+3.794032217540524)/2=3.14159=\pi.\/$
[…]
Orbit parameters under consideration are compared in the following table for the reader’s convenience:
Parameters: $U^{91+}$ Ion $Og^{117+}$ Ion $Ubb^{121+}$ Ion
$\omega$ 0.741 1350.741\ 135 0.508 4570.508\ 457 0.455 4190.455\ 419
$\epsilon$ 0.904 8820.904\ 882 0.941 4790.941\ 479 0.949 7820.949\ 782
$a/a_{0}$ 0.035 31630.035\ 3163 0.022 20410.022\ 2041 0.020 35340.020\ 3534
$r_{\text{min}}$ 0.003 3592 090.003\ 3592\ 09 0.001 29940.001\ 2994 0.001 022110.001\ 02211
$r_{\text{max}}$ 0.067 2734 720.067\ 2734\ 72 0.043 10870.043\ 1087 0.039 68470.039\ 6847
$\Delta\theta$ 2.194 612.194\ 61 6.074 186.074\ 18 7.513 37.513\ 3
‘Winding’ number 11 22 33
Further animation details can be found in our complementary Mathematica notebooks [4].
3. Conclusion
We have demonstrated that in a super strong static electric field, similar to the hypothetical Oganesson hydrogen-like ion, there is a possibility of self-intersecting trajectory, which is different from the standard classical orbits. A possible rotation of an incoming electron around the Oganesson nucleus during the scattering should also be investigated [29]. It would be interesting to analyse a possibility of a similar ‘Oganesson-type’ effect in the theory of gravitation [18] and to study non-standard consequences in quantum electrodynamics in the strong, U, versus super strong, Og, Coulomb fields [1, 9, 10, 12, 13, 14, 19, 22, 23, 27, 28]. For example, in the creation of an electron-positron pair, the electron may be captured by the nucleus directly from the internal orbit. Further nuclear reaction with one of Oganesson’s proton may result in neutron and neutrino thus contributing into the instability of this super heavy element. In a hypothetical Unbibium, element with Z=122,Z=122\/, there is a possibility of a multiple self-intersecting trajectory, which should even more contribute into an orbit electron capture and the corresponding Ubb transformation.
Acknowledgments We are grateful to Dr. Ruben Abagyan, Dr. Sergey I. Kryuchkov, Dr. Eugene Stepanov, and Dr. Alexei Zhedanov for their insightful comments and help.